# Diffusion Codes: Self-Correction from Small(er)-Set Expansion with Tunable Non-locality

### v1

@@ 1 Introduction
What makes all these codes ‘good’ is expansion: expander graphs have large boundary-to-volume ratios and small diameters, properties that translate into constant rate and large distance for LDPC constructions. The catch is that such codes are not geometrically local: their parity checks couple qubits that are far apart in any low-dimensional Euclidean embedding. By contrast, in any fixed spatial dimension with local checks there are rigorous tradeoffs that preclude ‘good’ quantum LDPC codes and severely constrain mechanisms for passive memory [BPT10, BT09, Yos11]. This has sparked interest in a middle ground: physically realizable architectures that preserve some expander-like connectivity – such as codes with sparse long-range links, or fractal geometries. Exploring this middle regime is compelling both practically (lower overhead, higher thresholds, and potentially larger energy barriers) and conceptually, as it interpolates between the rigorously understood, effectively infinite-dimensional limit and the physically realized local Euclidean setting. This has motivated a program of study towards this effort [YC25, BDS+25, BG24, DL24, WB24, BW24].
@@
Against this backdrop, we develop a concrete “middle-ground” construction that interpolates between fully local Euclidean lattices and fully nonlocal expanders. We introduce a family of codes which we call diffusion codes, with a controllable knob that trades locality for expansion, yielding classical (and, via hypergraph product, quantum) LDPC codes with, what we term, ”smaller set expansion.” In typical expander codes, errors up to a size which is extensive in the total number of bits trigger a syndrome which grows at least linearly with the size of the error. This is known in the literature as ”small set expansion.” In contrast, our diffusion codes are smaller set expanders, which means only errors up to a size which is sub-extensive in system size have a linearly growing (in error weight) lower bound on the syndrome weight. We define these terms in more detail in Section 2.
@@
The construction of these codes arises from a local shuffling process on the Tanner graph which resembles a diffusion process, whence the name. We rigorously prove our results on the cycle graph which realizes a quasi-1D classical code (quasi-2D quantum code), but we expect the results to generalize to arbitrary graphs.
@@
Diffusion codes have bits (qubits) which are arranged on a finite dimensional Euclidean lattice. The models are non-local, but the checks of the code involve bits which are all contained within a tunable radius of size $\delta(n)\sim n^{\beta}$, which scales sub-extensively with system size ($\beta<1$). Within this radius, the interaction graph of the code resembles that of an expander code; globally, the Euclidean structure reemerges. We prove that even for arbitrarily small $\beta>0$, the code retains small(er) set expansion: subsets of vertices up to size $\delta(n)$ are guaranteed to have large boundaries. Classical diffusion codes have parameters scaling as $n,\mathcal{O}(n),\mathcal{O}(n^{\beta})]$. Importantly, the smaller-set expansion property is still sufficient to prove that our codes are self-correcting. We show that the quantum codes obtained via hypergraph product inherit the same smaller-set expansion and self-correction properties. We conclude by showcasing some numerical experiments that illustrate the properties that we prove.
@@
We prove the above results by relating expansion in diffusion codes to properties of the joint distribution of interparticle-spacings in a simple exclusion process induced by the SWAP network which corresponded to the aforementioned shuffling process. The key technical challenge is to establish properties of this distribution for arbitrary initial states and at “mesoscopic” times $T$, which are longer than the subsystem size $\delta$ but much shorter than the global mixing time ($\delta\ll T\ll t_{\rm mix}$). We expect our proof technique for achieving this to be of independent interest.
@@ 1.1 Construction of Diffusion Codes
Smaller-set expansion of the Tanner graph has many direct consequences for the properties of the resulting code. For $\gamma>w_{\rm bit}/2$ (or equivalently, $\varepsilon<1/2$), it implies codes expansion up to error size $\delta(n)$ in the sense of Sipser and Spielman (Definition 2.18). This in turn implies that as long as $\delta(n)$ scales as an arbitrarily small power law $\delta\sim n^{\beta}$ for some $\beta>0$, the codes are self correcting (see Section 3.3). For $\gamma>3w_{\rm bit}/4$ [LTZ15b] it results in expansion up to error size $\delta(n)$ in the hypergraph product of the code with itself. Again, as long as $\delta\sim n^{\beta}$ for some $\beta<1$, this implies quantum self correction.
@@ 1.4 Related Work
Here, we give a brief overview of related work.
@@
Diffusion codes in 1D constructed from the cycle graph are in some ways similar to this idea. The size of the protograph is analagous to the scale of expansion. In fact, consider a diffusion code of size $n$ constructed from a diffusion time of $T\sim n^{2\beta}$. This code has code parameters $n,\mathcal{O}(n),\mathcal{O}(n^{\beta})]$. These same code parameters could be achieved by placing $n^{1-\beta}$ different ”good” codes of size $n^{\beta}$ along a line, even without coupling them. What separates diffusion codes construction from these is that their construction is stochastically homogeneous, making them more natural from a physics perspective, and also potentially easier to implement.
@@
Embedding from higher dimensions. Recently, there have been a few examples of codes saturating the BPT bounds ([BPT10]) in finite dimensions. These codes are all constructed via algebraic embeddings of higher dimensional optimal codes into lower dimensions while sacrificing as little favorable properties of the input codes as possible[WB24, Por23, BW24]. However, in general at least some favorable properties of the input codes will be lost. For example, it has recently been shown that layer codes are not self-correcting, even if the input codes are [Bas25].
@@
Long range models. When viewed as Hamiltonians, the diffusion codes we construct, especially from the cycle graph, can be viewed as long-range models, where the scale of any interaction has spread over a power law in system size. This construction takes some inspiration from a vast body of literature in physics concering long-range models. Indeed a famous example is of the 1D Ising model. Though it normally has no phase transition, if one instead considers a 1D Ising model where every spin can interact with every other spin with a strength that falls off as a power law in distance, then, depending on the strength of the power law, the resulting model does have a phase transition [Dys69]. Another example is of a 1D embedding of the famous Sherrington-Kirkpatrick (SK) model of spin glasses [LPRTRL08]. In this model, the authors embed the SK model onto a 1D lattice and every interaction $J_{ij}$ has some probability of being zero which decays as a power law in |ij||i-j|. Through extensive numerical experiments, they present evidence of replica symmetry breaking and spin glass phenomenology.
@@ 5 Conclusion and Future Direction
    Fewer nonlocal checks: In the diffusion code construction, although non-locality can be turned, almost all edges are of a macroscopic length. This remains true when considering the quantum codes constructed from them. There have been some recent results by Li & Dai that discuss the degree of nonlocality needed for 2D quantum codes in order to surpass the BPT bounds [BPT10, BT09, DL24]. There, the claim is that any $[n,k,d]$ 2D quantum stabilizer code with $kd^{2}\geq\mathcal{O}(n)$ must have at least $c_{0}\cdot\max(k,d)$ interactions of length at least $c_{0}\cdot\max\left(\frac{d}{\sqrt{n}},\left(\frac{kd^{2}}{n}\right)^{1/4}\right)$. The diffusion codes may be tuned to meet the requisite length scale, but a simple extension of Lemma 3.38 will show that almost surely the length of every interaction will be of this length, not just the minimum number required. From this perspective, our quantum codes are sub-optimal. We speculate that their may be other stochastic processes or modifications to the interchange process (such as introducing a weak attractive interaction) which could lead to a more tailored edge-length distribution. In particular, it would be ideal to be able to tune the geometric size distribution of the checks, for example to have the distribution be exponential. Numerics suggest that this would enable fault-tolerant implementation of such codes even using 2D local circuits [BG24, BDS+25].
@@
    Practical Implementations: Part of the motivation for our work has also been the development of quantum simulator platforms and near term quantum devices [doi21, blu24]. These new experimental platforms have shown rapid development in recent years and have also been used for error correction experiments [goo25]. The long range connectivity of quantum expander codes makes it challenging to implement these on current platforms, though there is some intriguing work in this direction [CKPSS24, GHKL25]. Given that the degree of nonlocality may be tuned in the diffusion codes, we are curious whether they may be more easily implemented in current platforms such as neutral atom simulators. The construction of diffusion codes suggests a natural SWAP circuit for syndrome extraction, though the circuit depth is sub optimal [DBT21]. Speeding up the mixing time of the underlying shuffling process (e.g. via lifting) would however directly translate into shorter syndrome extraction circuits.

### v2

@@ 1 Introduction
What makes all these codes ‘good’ is expansion: expander graphs have large boundary-to-volume ratios and small diameters, properties that translate into constant rate and large distance for LDPC constructions. The catch is that such codes are not geometrically local: their parity checks couple qubits that are far apart in any low-dimensional Euclidean embedding. By contrast, in any fixed spatial dimension with local checks there are rigorous tradeoffs that preclude ‘good’ quantum LDPC codes and severely constrain mechanisms for passive memory [BPT10, BT09, Yos11]. This has sparked interest in a middle ground: physically realizable architectures that preserve some expander-like connectivity—such as codes with sparse long-range links, or fractal geometries. Exploring this middle regime is compelling both practically (lower overhead, higher thresholds, and potentially larger energy barriers) and conceptually, as it interpolates between the rigorously understood, effectively infinite-dimensional limit and the physically realized local Euclidean setting. This has motivated a program of study towards this effort [YC25, BDS+25, BG24, DL24, WB24, BW24].
@@
Against this backdrop, we develop a concrete “middle-ground” construction that interpolates between fully local Euclidean lattices and fully nonlocal expanders. We introduce a family of codes which we call diffusion codes, with a controllable knob that trades locality for expansion, yielding classical (and, via hypergraph product, quantum) LDPC codes with, what we term, “smaller set expansion”. In typical expander codes, errors up to a size which is extensive in the total number of bits trigger a syndrome which grows at least linearly with the size of the error. This is known in the literature as ”small set expansion.” In contrast, our diffusion codes are smaller set expanders, which means only errors up to a size which is sub-extensive in system size have a linearly growing (in error weight) lower bound on the syndrome weight. We define these terms in more detail in Section 2. We state this main result as an informal theorem below:
Theorem 1.1 (Main Result, Informal).
For all $\beta>0$, there exist classical $n,\mathcal{O}(n),\mathcal{O}(n^{\beta})]$ LDPC codes with $\beta>0$, and an embedding of these codes onto the line such that
  1. 1.
    the geometric size of every check (the diameter of its support on the line) is concentrated around $\sim\mathcal{O}(n^{\beta})$
  2. 2.
    the code has the property of “smaller set” expansion, i.e. there exist constants $(\gamma,\delta)$ for any error $\mathbf{x}$ such that $|\mathbf{x}|\leq\delta\cdot n^{\beta}$, the weight of the associated syndrome $\mathbf{s}$ is lower bounded as $\mathbf{s}\geq\gamma\cdot|\mathbf{x}|$.
By taking the hypergraph product of two such codes, one can further obtain a family of quantum LDPC codes with the same locality in two dimensions and the same “smaller set” expansion property.
@@
The construction of these codes arises from running a local shuffling process on the Tanner graph for a fixed amount of time. The process resembles diffusion, whence the name. We rigorously prove that this results in codes with smaller set expansion on the cycle graph, which yields a quasi-1D classical code (quasi-2D quantum code via the hypergraph product). However, based on the physically expected generality of diffusion ideas, we expect the results to actually hold on arbitrary graphs.
@@
Diffusion codes have bits (qubits) which are arranged on a finite dimensional Euclidean lattice. The models are non-local, but the checks of the code involve bits which are all contained within a tunable radius of size $\delta(n)\sim n^{\beta}$, which scales sub-extensively with system size ($\beta<1$). This radius is tuned by the time that the shuffling process is run. The result is that locally within this radius, the interaction graph of the code resembles that of an expander code; globally, the Euclidean structure reemerges. We prove that even for arbitrarily small $\beta>0$, the code retains small(er) set expansion: subsets of vertices up to size $\delta(n)$ have a boundary of a size proportional to the subset size. It follows that classical diffusion codes can have parameters scaling as $n,\mathcal{O}(n),\mathcal{O}(n^{\beta})]$. Importantly, the smaller-set expansion property is still sufficient to prove that our codes are self-correcting. We show that the quantum codes obtained as the hypergraph product of two classical diffusion codes inherit the same smaller-set expansion and self-correction properties. We conclude by showcasing some numerical experiments that illustrate the properties that we prove.
@@
We prove the above results by relating expansion in diffusion codes to properties of the joint distribution of interparticle spacings in a simple exclusion process induced by the aforementioned shuffling process. The key technical challenge is to establish properties of this distribution for arbitrary initial states and at “mesoscopic” times $T$, which are longer than the subsystem size $\delta$ but much shorter than the global mixing time ($\delta\ll T\ll t_{\rm mix}$). We expect our proof technique for achieving this to be of independent interest.
@@ 1.1 Construction of Diffusion Codes
Smaller-set expansion of the Tanner graph has many direct consequences for the properties of the resulting code. For $\gamma>w_{\rm bit}/2$ (or equivalently, $\varepsilon<1/2$), it implies code expansion up to error size $\delta(n)$ in the sense of Sipser and Spielman (Definition 2.18). This in turn implies that as long as $\delta(n)$ scales as an arbitrarily small power law $\delta\sim n^{\beta}$ for some $\beta>0$, the codes are self correcting (see Section 3.3). For $\gamma>3w_{\rm bit}/4$, [LTZ15b] smaller-set expansion implies code expansion up to error size $\delta(n)$ in the hypergraph product of the code with itself. Again, as long as $\delta\sim n^{\beta}$ for some $\beta<1$, this implies quantum self correction.
@@ 1.4 Related Work
Here, we give a brief overview and comparison of related work.
@@
Diffusion codes in 1D constructed from the cycle graph are in some ways similar to this idea. The size of the protograph is analagous to the scale of expansion. In fact, consider a diffusion code of size $n$ constructed from a diffusion time of $T\sim n^{2\beta}$. This code has code parameters $n,\mathcal{O}(n),\mathcal{O}(n^{\beta})]$. These same code parameters could be achieved by placing $n^{1-\beta}$ different ”good” codes of size $n^{\beta}$ along a line, even without coupling them. What sets diffusion codes apart from these is the simplicity and flexibility of their construction (they naturally can be defined on any graph geometry). Diffusion codes are also by construction stochastically homogeneous with respect to the underlying graph, making them more natural from a physics perspective, and also potentially easier to implement.
Interpolation from Topological Codes. Another approach to achieve tunable non-locality proposed in recent years has been to add non-locality to topological codes. This may be done either by modifying some number of checks to make them long range [HMKL24], or to insert additional checks and bits to “augment” the topological code [RWBC20]. Yet another approach is to concatenate an appropriate qLDPC code with surface codes, in which case the quasi-locality may be tuned by varying the input qLDPC code [DL24]. This last construction has optimal code parameter tradeoffs, in the sense that itsaturates generalized BPT bounds, which were presented in the same work. These generalized bounds answer the question of how many long range interactions in 2D are required in order to circumvent the BPT bound. Dai and Li proved that any $[n,k,d]$ 2D quantum stabilizer code with $kd^{2}\geq\mathcal{O}(n)$ must have at least $c_{0}\cdot\text{max}(k,d)$ interactions of length at least $c_{0}\cdot\text{max}\left(\frac{d}{\sqrt{n}},\left(\frac{kd^{2}}{n}\right)^{1/4}\right)$. By concatenating a good qLDPC code with a surface code and implementing the result on a “stacked architecture” [DL24] they are further able to interpolate the trade-off between the number of nonlocal checks, and their size.
These approaches are interesting, and allow for much flexibility. Our construction is quite different in nature, and in contrast to all approaches summarized above naturally homogeneous in space. In terms of parameters, in our construction we have many checks which will typically be of size $L^{\beta}$, where $\beta$ can be made arbitrarily small (at the cost of also decreasing the distance by the same amount). We emphasize that while out construction is sub-optimal from a pure code parameter perspective, we proof several properties of diffusion code beyond just parameter scaling, such as expansion, and as a consequence self-correction and single-shot decoding.
@@
Embedding from higher dimensions. Recently, there have been a few examples of codes saturating the BPT bounds ([BPT10]) in finite dimensions. These codes are all constructed via algebraic embeddings of higher dimensional optimal codes into lower dimensions while sacrificing as little favorable properties of the input codes as possible [WB24, Por23, BW24]. More generally, from theory of metric embeddings, there exist methods to embed graphs into Euclidean dimensions (e.g. Johnson-Lindenstrauss lemma) at the cost of distortion and some nonlocality [Woo25]. However, in general at least some favorable properties of the input codes will be lost. For example, it has recently been shown that layer codes are not self-correcting, even if the input codes are [Bas25]. Baspin [Bas23] further showed that the connectivity graph of any $n,\Theta(n),d]$ quantum code must contain expanders of size $\Omega(d)$ as subgraphs, and the union of these subgraphs must cover a finite fraction of the full connectivity graph. This is similar (although not identical) to the Tanner graph of the code being an $\Omega(d)$ “smaller set” expander. In particular, the results do not allow for the recovery of lossless bipartite expanding subgraphs, which is necessary to guarantee the expansion properties of the hypergraph product.
@@
Long range models. When viewed as Hamiltonians, the diffusion codes we construct, especially from the cycle graph, can be viewed as long-range models, where the scale of any interaction has spread over a power law in system size. This construction takes some inspiration from a vast body of literature in physics concerning long-range models. Indeed a famous example is that of the 1D Ising model. Although the short-range version has no phase transition, if one instead considers a 1D Ising model where spins interact with a strength that falls off as a power law in distance, then, depending on the strength of the power law, the resulting model does have a phase transition [Dys69]. Another example is a one-dimensional embedding of the Sherrington-Kirkpatrick (SK) model of spin glasses presented in [LPRTRL08]. The authors embed the SK model onto a 1D lattice by considering an interaction matrix $J_{ij}$, where each entry has a probability of being non-zero which decays as a power law in |ij||i-j|. Through extensive numerical experiments, they present evidence of replica symmetry breaking and spin glass phenomenology in this model.
@@ 5 Conclusion and Future Direction
    Fewer nonlocal checks: In the diffusion code construction, although non-locality can be turned, almost all edges are of a macroscopic length. This remains true when considering the quantum codes constructed from them. As mentioned earlier, recent results by Dai and Li discuss the degree of nonlocality needed for 2D quantum codes in order to surpass the BPT bounds [BPT10, BT09, DL24]. To reiterate, they prove that any $[n,k,d]$ 2D quantum stabilizer code with $kd^{2}\geq\mathcal{O}(n)$ must have at least $c_{0}\cdot\max(k,d)$ interactions of length at least $c_{0}\cdot\max\left(\frac{d}{\sqrt{n}},\left(\frac{kd^{2}}{n}\right)^{1/4}\right)$. The diffusion codes may be tuned to meet the requisite length scale, but a simple extension of Lemma 3.38 will show that almost surely the length of every interaction will be of this length, not just the minimum number required. From this perspective, our quantum codes are sub-optimal. We speculate that their may be other stochastic processes or modifications to the interchange process (such as introducing a weak attractive interaction) which could lead to a more tailored edge-length distribution. In particular, it would be ideal to be able to tune the geometric size distribution of the checks, for example to have the distribution be exponential. Numerics suggest that this would enable fault-tolerant implementation of such codes even using 2D local circuits [BG24, BDS+25].
@@
    Practical Implementations: Part of the motivation for our work has also been the development of quantum simulator platforms and near term quantum devices [doi21, blu24]. These new experimental platforms have shown rapid development in recent years and have also been used for error correction experiments [goo25]. The long range connectivity of quantum expander codes makes it challenging to implement these on current platforms, though there is some intriguing work in this direction [CKPSS24, GHKL25]. Given that the degree of nonlocality may be tuned in the diffusion codes, we are curious whether they may be more easily implemented in current platforms such as neutral atom simulators. The construction of diffusion codes also suggests a natural SWAP circuit for syndrome extraction, though the circuit depth is sub optimal [DBT21]. Speeding up the mixing time of the underlying shuffling process (e.g. via lifting) would however directly translate into shorter syndrome extraction circuits.